# Progressive-Proximity Bit-Flipping for Decoding Surface Codes

### v1

@@ IV-1 Efficient computation of the proximity vector
Instead, we exploit the quasi-cyclic property of the toric code to make the procedure more efficient. The idea is to pre-compute the proximity influence of an arbitrary check node, say $\boldsymbol{\gamma}(c_{1}),\boldsymbol{\nu}(c_{1})$, store it prior to the decoding process, and then compute online $\boldsymbol{\gamma}(c_{i}),\boldsymbol{\nu}(c_{i})$ when needed, by appropriately permuting $\boldsymbol{\gamma}(c_{1}),\boldsymbol{\nu}(c_{1})$. Assume that we have an $L\times L$ toric code. We can label the variable nodes with non-negative integers row-wise in increasing order (so that the first row is labeled from 00 to $L-1$, the second row is labeled from $L$ to $12L-$ 1, and so on), and we can do the same for the check nodes. An example is provided in Fig. 1 for the case $L=3$. There will be a total of $2L^{2}$ variable nodes and $L^{2}$ check nodes. Assume that we have already stored $\boldsymbol{\nu}(c_{i})$ and $\boldsymbol{\gamma}(c_{i})$ for some arbitrary check node $c_{i}$, and that we now wish to compute $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, for some $j\neq i$, by appropriately permuting $\boldsymbol{\nu}(c_{i})$ and $\boldsymbol{\gamma}(c_{i})$. By exploiting the indexing we have defined for the check and variable nodes, we can define a vertical shift $\sigma_{y}$ and an horizontal shift $\sigma_{x}$ in the following manner:
$\begin{cases}\sigma_{y}=\lfloor(j-i)/L\rfloor\\ \sigma_{x}=j-i\mod L\ \ ,\end{cases}$ (13)
@@
assuming that $i<j$. We can express the index transformation in a closed form, using $\sigma_{x}$ and $\sigma_{y}$. Let $k_{c}\in[1,L^{2}]$ and $k_{v}\in[1,2L^{2}]$ be the indices of the check and variable nodes, respectively. We can express a coordinate shift using two linear maps from $k_{c}$ to $k^{\prime}_{c}$ and from $k_{v}$ to $k^{\prime}_{v}$ respectively, such that
$\begin{cases}k^{\prime}_{c}=\bigl{\{}\{k_{c}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{c}-1)/L\rfloor+\sigma_{y}L\bigr{\}}\mod L^{2}\\ k^{\prime}_{v}=\bigl{\{}\{k_{v}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{v}-1)/L\rfloor+2\sigma_{y}L\bigr{\}}\mod 2L^{2}\\ \end{cases}$ (14)
Hence, we have $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=\gamma_{k_{c}}^{\ell}(c_{i})$ and $\nu_{k^{\prime}_{c}}^{\ell}(c_{j})=\nu_{k_{c}}^{\ell}(c_{i})$. An example of this transformation is illustrated in Fig. 2. In Fig 1(a), the proximity influence of $c_{4}$ of depth 1, namely $\boldsymbol{\nu}^{\ell}(c_{4})$ and $\boldsymbol{\gamma}^{\ell}(c_{4})$, is assumed to have been pre-computed. In Fig. 1(b), the influence is shifted to compute $\boldsymbol{\nu}^{\ell}(c_{8})$ and $\boldsymbol{\gamma}^{\ell}(c_{8})$; in the example, $\sigma_{x}=\sigma_{y}=1$. The reader can verify that, using (14), the variable node $4$ is mapped to $11$, variable node $6$ to $13$ and so on; similarly, the influence on the checks is shifted. We utilize the $\mathtt{Shift\_influence}$ subroutine to apply the coordinate permutation of (14); the subroutine, described in Algorithm 1, takes as input the index of a check node $c_{j}$, and applies (14) to $\boldsymbol{\gamma}(c_{1})$ and $\boldsymbol{\nu}(c_{1})$ to obtain $\boldsymbol{\gamma}(c_{j})$ and $\boldsymbol{\nu}(c_{j})$.
@@
Global: $\boldsymbol{\gamma}(c_{1})$, $\boldsymbol{\nu}(c_{1})$, $L$
@@
1:$\mathbf{k}_{v}\leftarrow[1,2L^{2}]$
$\mathbf{k}_{c}\leftarrow[1,L^{2}]$
3:Compute $\sigma_{x},\sigma_{y}$ using (13)
4:Compute $\mathbf{k}^{\prime}_{v},\mathbf{k}^{\prime}_{c}$ using (14)
5:$\boldsymbol{\nu}_{\mathbf{k}^{\prime}_{v}}(c_{j})\leftarrow\boldsymbol{\nu}_{\mathbf{k}_{v}}(c_{1})$
6:$\boldsymbol{\gamma}_{\mathbf{k}^{\prime}_{c}}(c_{j})\leftarrow\boldsymbol{\gamma}_{\mathbf{k}_{c}}(c_{1})$
7:return
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(a) The depth-1 influence of $c_{4}$ (highlighted in red) is computed; check and variable nodes such that $\nu(c_{4})\neq 0$ and $\gamma(c_{4})\neq 0$ are represented in gray, while nodes such that $\nu(c_{4})=0$ and $\gamma(c_{4})=0$ are represented in white.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(b) The influence of Fig. 1(a) is shifted to the node $c_{8}$: in this case, $\sigma_{x}=\sigma_{y}=1$.
Figure 2: Example of shifting of the proximity influence from $c_{5}$ to $c_{8}$.
@@ IV-2 Auxiliary proximity influence
We also define an auxiliary proximity influence of a check node $c_{j}$, that we denote as $\mathbf{v}(c_{j})$.
@@
The auxiliary proximity influence $\mathbf{v}(c_{j})$ is a $1\times n$ vector such that $v_{i}(c_{j})$ is the length of the shortest path between the variable node $v_{i}$ and the check node $c_{j}$. For example, if a variable node $v_{i}$ is directly connected to $c_{j}$, then $v_{i}(c_{j})=1$.
Note that the proximity influences $\boldsymbol{\nu}(c_{j})$ and $\mathbf{v}(c_{j})$ share the same non-zero positions, but in general they have different values: while $\boldsymbol{\nu}(c_{j})$ will have higher values for the variable nodes close to $c_{j}$, in $\mathbf{v}(c_{j})$ the variable nodes connected to $c_{j}$ will have value $1$, those at distance 2 will have value $2$, and so on. We can compute $\mathbf{v}(c_{j})$ in a similar way to $\boldsymbol{\nu}(c_{j})$. Thus, it is sufficient to construct $\mathbf{v}(c_{1})$ offline, and apply Algorithm 1 to create $\mathbf{v}(c_{j})$, for any $j$. We illustrate in Section V-A1 how we make use of the auxiliary proximity influence to correct errors. Finally, we define two subroutines, illustrated in Algorithm 2 and Algorithm 3. The $\mathtt{Compute\_proximity\_vector}$ subroutine takes as input the syndrome $\mathbf{s}$ and generates the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, using Algorithm 1 to compute individual influences and summing them up according to (11). The $\mathtt{Shift\_and\_remove}$ subroutine takes as input the residual syndrome $\mathbf{s}^{\prime}$, as well as the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, and updates them using (12).
@@ V-A Preliminary BF
V-A1 The iterative-matching routine
[…]
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
@@ V-A1 The iterative-matching routine
(a) Proximity influence of depth 2 for the check node $c_{8}$; different coloring of variable nodes correspond to their distance from $c_{8}$: orange nodes are at distance 1, yellow at distance 2.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(b) Proximity influence of depth 2 for $c_{4}$.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
@@
(c) Combination of the previous two proximity influences: variable nodes which were assigned yellow and orange are now colored red; those nodes which were assigned yellow/yellow are now deep yellow. Finally, the variable nodes which were assigned orange/white or yellow/white have maintained their color. The set of red variable nodes coincide with the union of all the possible error patterns that satisfy the matching, i.e., $\{v_{10},v_{13}\}$ and $\{v_{7},v_{11}\}$.
@@
Figure 3: Example of the usage of the auxiliary proximity influence of checks in Algorithm 6. In the example, $c_{4}$ and $c_{8}$ are unsatisfied checks, and the decoder has to find the shortest path between them, assuming to start with $c_{8}$ as pivot.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
@@
Figure 4: Addition of the extra check as described in Section V-A1. The blue variable nodes are the depth-1 influence of the extra check, and the darker ones are those which will be flipped. In the example, $\Delta x=\Delta y=1$.
@@
7:     $\mathbf{c}(c_{i}),\mathbf{v}(c_{i})\leftarrow\mathtt{Shift\_influence}(i)$
@@
10:     $\mathbf{c}(c_{j}),\mathbf{v}(c_{j})\leftarrow\mathtt{Shift\_influence}(j)$
@@
16:         if $j-i\mod L\ <i-j\mod L$ then
17:              $\Delta x\leftarrow j-i\mod L$
18:         else
19:              $\Delta x\leftarrow-(i-j\mod L)$
20:         end if
21:         $\Delta y\leftarrow\mathrm{min}\{\lfloor\frac{i-1}{L}\rfloor-\lfloor\frac{j-1}{L}\rfloor\mod L,\ \lfloor\frac{j-1}{L}\rfloor-\lfloor\frac{i-1}{L}\rfloor\mod L\}$
22:         $z\leftarrow i+\Delta x\mod L+L\lfloor\frac{i-1}{L}\rfloor$
23:         $\mathbf{c}(c_{z}),\mathbf{v}(c_{z})\leftarrow\mathtt{Shift\_influence}(z)$
24:         $\mathbf{f}_{1}\leftarrow k\ |\ v_{k}(c_{i})+v_{k}(c_{z})=\Delta x+1$
25:         $\mathbf{f}_{2}\leftarrow k\ |\ v_{k}(c_{j})+v_{k}(c_{z})=\Delta y+1$
26:         $\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\leftarrow\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\oplus 1$
27:     end if
28:     $\mathbf{s}^{\prime}\leftarrow\hat{\mathbf{e}}\cdot\mathbf{H}^{T}\mod{2}$
29:     $\boldsymbol{\gamma}^{\prime}\leftarrow\mathtt{Shift\_and\_remove}(\cdot,\boldsymbol{\gamma},\mathbf{s}^{\prime})$
30:     $\hat{\mathbf{s}}\leftarrow\mathbf{s}\oplus\mathbf{s}^{\prime}$
31:end while
32:return $\hat{\mathbf{e}}$
V-A2 Decoding of non-toric codes
It is possible to generalize the algorithm for non-toric surface codes, i.e., for those codes which present open borders, such that the variable nodes at the boundaries will have degree 1. An example that we present in this paper is the rotated planar code [28]; the X-Tanner graph of the $\llbracket 9,1,3\rrbracket$ rotated planar code is illustrated in Fig. 5.
In principle, our decoder is not suitable for non-toric codes, due to degree-1 variable nodes: indeed, if one of those variable nodes is in error, they can not be corrected by the BF because they only have degree 1, and neither by the iterative matching procedure, because there will only be one unsatisfied check, while we need at least two to perform a matching. Note that every matching decoder for a surface code needs to deal with this problem. To solve this issue, we adopt a similar solution solution to that proposed for MWPM in [2]: we attach virtual checks to the boundaries of the code, as shown in Fig. 5. The virtual checks are only considered in the iterative matching procedure; they are always considered to be unsatisfied, and the matching decoder treats them as normal checks, except for the fact that a virtual check cannot be chosen as a starting pivot.
00001111222211113333444455552222666677778888333344445555
Figure 5: Tanner graph for the $\llbracket 9,1,3\rrbracket$ rotated planar code. Black check nodes with dashed edges represent virtual check nodes.
Note that now, with open boundaries, we can still use Algorithms 1 and 2, but (14) has to be reformulated to take into account the open boundaries:
$\begin{cases}k^{\prime}_{c}=(k_{c}-\sigma_{x}+\lfloor(k_{c}-1)/L\rfloor-\sigma_{y}L)\ \mathrm{if}\ k^{\prime}_{c}\leq L^{2}\\ k^{\prime}_{v}=(k_{v}-\sigma_{x}+\lfloor(k_{v}-1)/L\rfloor-2\sigma_{y}L)\ \mathrm{if}\ k^{\prime}_{v}\leq 2L^{2}\\ k^{\prime}_{c},k^{\prime}_{v}=-1\ \mathrm{otherwise.}\end{cases}$ (15)
In this way, the proximity influence does not extend out of the graph. Also, for virtual check nodes we assign arbitrary proximity values; in other words, $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=\gamma_{k_{c}}^{\ell}(c_{i})$ if $c_{i}$ is not an auxiliary check and $k_{c}^{\prime}\leq L^{2}$, $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=0$ if $c_{i}$ is not an auxiliary check and $k_{c}^{\prime}=-1$, and $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=C$, where $C$ is an arbitrary constant value, if $c_{i}$ is an auxiliary check. The rest of the algorithm remains the same.

### v2

@@ IV-1 Efficient computation of the proximity vector
Instead, we exploit the labeling we have assigned to variable and check nodes in the surface code. The idea is to pre-compute the proximity influence of an arbitrary check node over a larger surface code, say $\boldsymbol{\gamma}(c_{1}),\boldsymbol{\nu}(c_{1})$, store it prior to the decoding process, and then compute online $\boldsymbol{\gamma}(c_{i}),\boldsymbol{\nu}(c_{i})$ when needed by appropriately permuting $\boldsymbol{\gamma}(c_{1})$ and $\boldsymbol{\nu}(c_{1})$. We illustrate this process for the rotated surface code, although it can be extended for other types of surface codes. Assume that we label the variable nodes with non-negative integers row-wise in increasing order and we do the same for the check nodes; an example is provided in Fig. 1 for a rotated code with $L=3$. Consider the Tanner graphs $\mathcal{G}_{1}$ of a rotated code with $L=L_{1}$, and $\mathcal{G}_{2}$ of an $L_{2}^{\prime}\times L_{2}^{\prime\prime}$ surface code , such that $L_{2}>L_{1}$ and $L_{2}^{\prime\prime}>L_{1}$. Let $\mathcal{L}^{(1)}_{v},\mathcal{L}^{(1)}_{c}\in\mathbb{N}^{L_{1}^{2}}$ be the labels of the variable and check nodes of $\mathcal{G}_{1}$, and $\mathcal{L}^{(2)}_{v},\mathcal{L}^{(2)}_{c}\in\mathbb{N}^{L_{2}^{2}}$ be the labels of the variable and check nodes of $\mathcal{G}_{2}$, respectively. There are two injective mappings $\phi_{c}:\mathcal{L}^{(1)}_{c}\rightarrow\mathcal{L}^{(2)}_{c}$ and $\phi_{v}:\mathcal{L}^{(1)}_{v}\rightarrow\mathcal{L}^{(2)}_{v}$ from the labels of the check and variable nodes of $\mathcal{G}_{1}$ to the labels of the check and variable nodes of $\mathcal{G}_{2}$. Let $i,j\in\mathcal{L}^{(1)}_{c}$. Assume that we have calculated $\boldsymbol{\nu}(c_{\phi_{c}(i)})$ and $\boldsymbol{\gamma}(c_{\phi_{c}(i)})$ for some arbitrary check node $c_{i}$ using (10) on $\mathcal{G}_{2}$, and that we now wish to compute $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, for some $j\neq i$, by appropriately permuting $\boldsymbol{\nu}(c_{\phi_{c}(i)})$ and $\boldsymbol{\gamma}(c_{\phi_{c}(i)})$. Let $\rho_{c}$ be the number of check in each row of $\mathcal{G}_{2}$ and $\rho_{vo},\rho_{ve}$ be the number of variable nodes in even and odd rows of $\mathcal{G}_{2}$, respectively. By exploiting the indexing we have defined for the check and variable nodes, we can define a vertical shift $\sigma_{y}$ and an horizontal shift $\sigma_{x}$:
$\begin{cases}\sigma_{y}=\lfloor\frac{\phi_{c}(j)}{\rho_{c}}\rfloor-\lfloor\frac{\phi_{c}(i)}{\rho_{c}}\rfloor,\\ \sigma_{x}=\phi_{c}(j)\mod\rho_{c}\ -\ \phi_{c}(i)\mod\rho_{c}\ ,\\ \end{cases}$ (14)
@@
Note that $\sigma_{x}$ and $\sigma_{y}$ can also assume negative values. Depending on the surface code we are considering, it is possible to express the index transformation in closed form, using $\sigma_{x}$ and $\sigma_{y}$. Let $k_{c},k^{\prime}_{c}\in\mathcal{L}_{c}^{(2)}$ and $k_{v},k^{\prime}_{v}\in\mathcal{L}_{v}^{(2)}$. We can express a coordinate shift using two linear maps from $k_{c}$ to $k^{\prime}_{c}$ and from $k_{v}$ to $k^{\prime}_{v}$ respectively, such that
$\begin{cases}k^{\prime}_{c}=k_{c}+\sigma_{x}+\sigma_{y}\rho_{c}\\ k^{\prime}_{v}=k_{v}+\sigma_{x}+\sigma_{y}(\rho_{vo}+\rho_{ve}).\end{cases}$ (15)
Thus, we have $\nu_{k^{\prime}_{v}}(c_{\phi_{c}(j)})=\nu_{k_{v}}(c_{\phi_{c}(i)})$ and $\gamma_{k^{\prime}_{c}}(c_{\phi_{c}(j)})=\gamma_{k_{c}}(c_{\phi_{c}(i)})$, for all $k^{\prime}_{c}$ and $k^{\prime}_{v}$; finally, we apply the inverse mappings $\phi_{c}^{-1}$, $\phi_{v}^{-1}$ to all the elements in the codomain of these functions, such that $\nu_{q}(c_{j})=\nu_{\phi_{v}(q)}(c_{\phi_{c}(j)})$ and $\gamma_{p}(c_{j})=\gamma_{\phi_{c}(p)}(c_{\phi_{c}(j)})$ for any $q\in\mathcal{L}_{v}^{(1)}$ and $p\in\mathcal{L}_{c}^{(1)}$. This procedure is illustrated in Algorithm 1, where $j\in\mathcal{L}^{(1)}_{c}$ is the label of the check nodes of which we want to compute the proximity vectors. An example of the application of Algorithm 1 is illustrated in Fig. 2. Specifically, we pick $\mathcal{G}_{1}$ to be a $L_{1}=3$ rotated code and $\mathcal{G}_{2}$ to be a $4\times 5$ surface code. The example illustrates how to shift the proximity vectors from $c_{\phi_{c}(0)=8}$ to $c_{\phi_{c}(2)=15}$. In Fig. 2(a) the proximity metrics $\boldsymbol{\gamma}(c_{\phi_{c}(0)})$ and $\boldsymbol{\nu}(c_{\phi_{c}(0)})$ are assumed to be pre-computed; in Fig. 2(b) we want to compute $\boldsymbol{\gamma}(c_{2})$ and $\boldsymbol{\nu}(c_{2})$, which is mapped to $\boldsymbol{\gamma}(c_{\phi_{c}(2)=15})$ and $\boldsymbol{\nu}(c_{\phi_{c}(2)=15})$. We compute $\sigma_{x}=1$ and $\sigma_{y}=1$ and apply (15) to each variable and check node. The reader can verify that, for instance, check node 7 is mapped to check node 14, and thus to check node $\phi^{-1}_{c}(14)=1$ in the $L=3$ surface code. On the other hand, check node 2 is mapped to check node 9, which doesn’t belong to the image of $\phi_{c}$, and thus $\boldsymbol{\gamma}(c_{\phi^{-1}(c_{9})})=\emptyset$. The values for $L_{2}^{\prime},L_{2}^{\prime\prime}$ are such that the mapping in (14) gives the correct result for every $\sigma_{x},\sigma_{y}$, which happens for $L_{2}^{\prime},L_{2}^{\prime\prime}>L+D$.
@@
Global: $\boldsymbol{\gamma}(c_{\phi_{c}(1)})$, $\boldsymbol{\nu}(c_{\phi_{c}(1)})$, $L_{2}$, $\phi_{v}$, $\phi_{c}$
@@
1:$\mathbf{k}_{v},\mathbf{k}_{c}\leftarrow[1,L_{2}^{2}]$
2:Compute $\sigma_{x},\sigma_{y}$ using (14)
3:Compute $\mathbf{k}^{\prime}_{v},\mathbf{k}^{\prime}_{c}$ using (15)
4:$\boldsymbol{\nu}_{\phi_{v}^{-1}(\mathbf{k}^{\prime}_{v})}(c_{j})\leftarrow\boldsymbol{\nu}_{\mathbf{k}_{v}}(c_{\phi_{c}(1)})$
5:$\boldsymbol{\gamma}_{\phi_{c}^{-1}(\mathbf{k}^{\prime}_{c})}(c_{j})\leftarrow\boldsymbol{\gamma}_{\mathbf{k}_{c}}(c_{\phi_{c}(1)})$
6:return
[…]
(a) The depth-1 influence of $c_{8}$ (highlighted in red) is computed on a $4\times 5$ surface code. The node label of $c_{8}$ corresponds to the node $c_{0}$ in the $L=3$ rotated code.
[…]
(b) The influence of Fig. 2(a) is shifted to the node $c_{15}$ on the $4\times 5$ surface code: in this case, $\sigma_{x}=1$ and $\sigma_{y}=1$. Then, the values of the shifted metrics are assigned to the nodes of the $L=3$ rotated code according to the illustrated mapping between labels.
Figure 2: Example of shifting of the proximity influence. The labeling on the variable (black) and check (red) nodes inside the parentheses is the labeling for the $L=3$ rotated code, while the labeling outside the parentheses is the one for the $4\times 5$ surface code. In this example, we compute the influence of $c_{15(2)}$ by shifting the influence of $c_{8(0)}$. The nodes colored in gray are the ones with a non-zero proximity metric; notice how some of the values are discarded after shifting, i.e. the values of $c_{9}$ and $c_{16}$.
@@ IV-2 Auxiliary proximity influence
We also define an auxiliary proximity influence of a check node $c_{j}$, that we denote as $\boldsymbol{\alpha}(c_{j})$.
@@
The auxiliary proximity influence $\boldsymbol{\alpha}(c_{j})$ is a $1\times n$ vector such that $\alpha_{i}(c_{j})$ is the length of the shortest path between the variable node $v_{i}$ and the check node $c_{j}$. For example, if a variable node $v_{i}$ is directly connected to $c_{j}$, then $\alpha_{i}(c_{j})=1$.
Note that the proximity influences $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\alpha}(c_{j})$ share the same non-zero positions, but in general they have different values: while $\boldsymbol{\nu}(c_{j})$ will have higher values for the variable nodes close to $c_{j}$, in $\boldsymbol{\alpha}(c_{j})$ the variable nodes connected to $c_{j}$ will have value $1$, those at distance 2 will have value $2$, and so on. We can compute $\boldsymbol{\alpha}(c_{j})$ in a similar way to $\boldsymbol{\nu}(c_{j})$. Thus, it is sufficient to construct $\boldsymbol{\alpha}(c_{1})$ offline, and apply Algorithm 1 to create $\boldsymbol{\alpha}(c_{j})$, for any $j$. We illustrate in Section V-B how we make use of the auxiliary proximity influence to correct errors. Finally, we define two subroutines, illustrated in Algorithm 2 and Algorithm 3. The $\mathtt{Compute\_proximity\_vector}$ subroutine takes as input the syndrome $\mathbf{s}$ and generates the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, using Algorithm 1 to compute individual influences and summing them up according to (11). The $\mathtt{Shift\_and\_remove}$ subroutine takes as input the residual syndrome $\mathbf{s}^{\prime}$, as well as the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, and updates them using (13).
@@ V-A Preliminary BF
V-B The iterative-matching routine
Here we present the iterative matching routine, which is specified in Algorithm 6. A matching decoder is a decoding algorithm specifically tailored for surface codes, which aims to pair together couples of unsatisfied checks, estimating the qubits on the shortest path connecting them as being in error. This comes from the fact that every error on the surface code can be interpreted as a path on the lattice, and the two endpoints of this path are the two unsatisfied checks generated by the error; however, errors occurring on boundary qubits only generate one unsatisfied check. To address this issue, it is common to add additional “dummy” unsatisfied check nodes to the boundaries of the surface code, as proposed in [2] (notice that this is not necessary for the toric code). These checks are illustrated in black in Fig. 3. We utilize the proximity vector $\boldsymbol{\gamma}$ to match pairs of unsatisfied checks together; specifically, we start by identifying the unsatisfied check $c_{i}$ with the lowest proximity vector entry $\gamma_{i}$ (line 4) calling it a pivot node, and compute its auxiliary proximity influence $\boldsymbol{\alpha}(c_{i})$. Among all the other unsatisfied checks, we pick the one at the smallest distance from the pivot (if there is more than one candidate at the same distance, we choose the check $c_{j}$ with lowest proximity vector entry $\gamma_{j}$); we call this the target node (line 9), and denote its distance from the pivot by $\delta$. After computing the auxiliary vector $\boldsymbol{\alpha}(c_{j})$, we compute $\boldsymbol{\alpha}(c_{i})+\boldsymbol{\alpha}(c_{j})$; the result of this operation is a vector such that if the $k$-th element is equal to $\delta+1$, the variable node $v_{k}$ belongs to the shortest path between $c_{i}$ and $c_{j}$. If the number of entries of $\boldsymbol{\alpha}(c_{i})+\boldsymbol{\alpha}(c_{j})$ that are equal to $\delta+1$ is equal to $\delta$, it means that there is only one shortest path between $c_{i}$ and $c_{j}$, that corresponds to the most likely error matching that syndrome; on the other hand, if the number of entries of $\boldsymbol{\alpha}(c_{i})+\boldsymbol{\alpha}(c_{j})$ that are equal to $\delta+1$ is larger than $\delta$, it means that the corresponding error is degenerate, as there is more than one shortest path between $c_{i}$ and $c_{j}$. An example of this is illustrated in Fig. 3, where there are two possible paths of length 2 between $c_{3}$ and $c_{7}$, and the number of variable nodes such that $\boldsymbol{\alpha}(c_{3})+\boldsymbol{\alpha}(c_{7})=3$ is more than 2. In the first case, it is sufficient to flip all the variable nodes associated with the value $\delta+1$, while in the latter case one of the possible degenerate errors must be chosen. We distinguish these two scenarios in line 13. To pick one among the possible paths, we compute the auxiliary influence of a third check node $c_{z}$, of position such that it is horizontally aligned with $c_{i}$ and vertically aligned with $c_{j}$; lines 16-22 are dedicated to computing the position $z$. In particular, we define $\sigma_{x}$ to be the horizontal shift between $c_{i}$ and $c_{j}$ and $\sigma_{y}$ to be the vertical shift between $c_{i}$ and $c_{j}$. Once $\boldsymbol{\alpha}(c_{z})$ is computed in lines 22-23, it is easy to check that the variable nodes $v_{k}$ such that $\alpha_{k}(c_{i})+\alpha_{k}(c_{z})=\sigma_{x}+1$ are those connecting $c_{i}$ and $c_{z}$, and those such that $\alpha_{k}(c_{j})+\alpha_{k}(c_{z})=\sigma_{y}+1$ are those connecting $c_{j}$ and $c_{z}$, thus we can flip all of them simultaneously in line 26.
00111100111122223333444422223333555566667777888899994444555510101010111111111212121213131313141414146666777715151515161616161717171718181818191919198888999920202020212121212222222223232323242424241010101011111111
@@ V-B The iterative-matching routine
(a) Proximity influence of depth 2 for the check node $c_{7}$; different colorings of variable nodes correspond to their distance from $c_{7}$: orange nodes are at distance 1, yellow at distance 2.
00111100111122223333444422223333555566667777888899994444555510101010111111111212121213131313141414146666777715151515161616161717171718181818191919198888999920202020212121212222222223232323242424241010101011111111
(b) Proximity influence of depth 2 for $c_{3}$.
00111100111122223333444422223333555566667777888899994444555510101010111111111212121213131313141414146666777715151515161616161717171718181818191919198888999920202020212121212222222223232323242424241010101011111111
@@
(c) Combination of the previous two proximity influences: variable nodes which were assigned yellow and orange are now colored red; those nodes which were assigned yellow and yellow are now deep yellow. Finally, the variable nodes which were assigned orange/white or yellow/white have maintained their color. The set of red variable nodes coincide with the union of all the possible error patterns that satisfy the matching, i.e., $\{v_{7},v_{12}\}$ and $\{v_{8},v_{13}\}$.
@@
Figure 3: Example of the usage of the auxiliary proximity influence of checks in Algorithm 6. In the example, $c_{3}$ and $c_{7}$ are unsatisfied checks, and the decoder has to find the shortest path between them, assuming it starts with $c_{7}$ as pivot.
00111100111122223333444422223333555566667777888899994444555510101010111111111212121213131313141414146666777715151515161616161717171718181818191919198888999920202020212121212222222223232323242424241010101011111111
@@
Figure 4: Addition of the extra check as described in Section V-B. The blue variable nodes are the depth-1 influence of the extra check, and the darker ones are those which will be flipped. In the example, $\sigma_{x}=\sigma_{y}=1$.
@@
7:     $\boldsymbol{\alpha}(c_{i})\leftarrow\mathtt{Shift\_influence}(i)$
@@
10:     $\boldsymbol{\alpha}(c_{j})\leftarrow\mathtt{Shift\_influence}(j)$
@@
16:         Get $\sigma_{x}$ using (14)
17:         Get $\sigma_{y}$ using (14)
18:         $z\leftarrow\phi_{c}^{-1}(\phi_{c}(z)+\sigma_{x}\rho_{c})$
19:         $\boldsymbol{\alpha}(c_{z})\leftarrow\mathtt{Shift\_influence}(z)$
20:         $\mathbf{f}_{1}\leftarrow k\ |\ v_{k}(c_{i})+v_{k}(c_{z})=\Delta x+1$
21:         $\mathbf{f}_{2}\leftarrow k\ |\ v_{k}(c_{j})+v_{k}(c_{z})=\Delta y+1$
22:         $\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\leftarrow\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\oplus 1$
23:     end if
24:     $\mathbf{s}^{\prime}\leftarrow\hat{\mathbf{e}}\cdot\mathbf{H}^{T}\mod{2}$
25:     $\boldsymbol{\gamma}^{\prime}\leftarrow\mathtt{Shift\_and\_remove}(\cdot,\boldsymbol{\gamma},\mathbf{s}^{\prime})$
26:     $\hat{\mathbf{s}}\leftarrow\mathbf{s}\oplus\mathbf{s}^{\prime}$
27:end while
28:return $\hat{\mathbf{e}}$
V-C Decoding of toric codes
In the case of toric codes it is possible to simplify the computation of the proximity vectors. Consider a $\llbracket 2L^{2},2,L\rrbracket$ toric code. The proximity vector of an arbitrary check node is pre-computed on the same code (there is no need to use a larger code), therefore for each node label $i$ we have $\phi(i)=i$; to perform the permutation from a node $i$ to a node $j$, $\sigma_{x}$ and $\sigma_{y}$ are computed using (14), and for each $k_{c},k_{v},k^{\prime}_{c},k^{\prime}_{v}\in[1,2L^{2}]$ we apply
$\begin{cases}k^{\prime}_{c}=\bigl{\{}\{k_{c}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{c}-1)/L\rfloor+\sigma_{y}L\bigr{\}}\mod L^{2}\\ k^{\prime}_{v}=\bigl{\{}\{k_{v}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{v}-1)/L\rfloor+2\sigma_{y}L\bigr{\}}\mod 2L^{2}\\ \end{cases}.$ (16)
To explain this procedure, we note that the toric code is composed of $L$ rows of $L$ check nodes, for a total of $L^{2}$ checks, and $2L$ rows of $L$ variable nodes, for a total of $2L^{2}$ variable nodes. The labels in each row are shifted cyclically modulo $L$ by the term $k_{c}+\sigma_{x}\mod L$, which returns a value in $[0,L-1]$. Adding $L\lfloor(k_{c}-1)/L\rfloor$ returns a value in the same row of $k_{c}$ but shifted by $\sigma_{x}$. Finally, a shift of $\sigma_{y}L$ modulo $L^{2}$ is applied to get the final label which will be in the same column but a different row. A similar reasoning is applied for $k_{v}$. The rest of the algorithm remains the same.